OpenStudy (anonymous):

Can someone help with Quadratic Functions? Find the solutions to the system y=x^2+8x+2 y=7x+4 I know first to set them and then from there but my difficulty is with expanding them. Once I get to there I get stuck. Please Help.

OpenStudy (anonymous):

@CookieCake

OpenStudy (anonymous):

Okay, for the first I got (-4,-14) How? Y=X^2+8X+2 X^2+8X+2=0 A=1 B=8 C=2 -b/2a (just plug the A and B into the formula.) -b/2a= -8/2(1) = -8/2= -4/1=-4 Axis of symmetry and X axis of vertex. X=-4 C-b^2/4a (2)-8^2/4(1) = 2-64/4= 2-16= -14 Y=-14 (-4,-14)

OpenStudy (anonymous):

The correct answer is (-2,-10) & (1,11) the thing is I dont know how and I know this is right for a fact.

OpenStudy (anonymous):

I'm sorry if my work did not help. I recommend a graphing calculator, or online math solver to check your work. Best of wishes! Again,sorry.

OpenStudy (anonymous):

Its fine & I have tried a solver and it cant help me I really dont get it & generally i'm alright with this kind of stuff. Thank you for trying, thats alot and I am really gracious for the time you did put into it.

OpenStudy (anonymous):

\[x^2+8x+2 =7x+4 \] right?

OpenStudy (anonymous):

set equal to zero and get \[x^2+x-2=0\] then factor as \[(x+2)(x-1)=0\] and now you can solve for \(x\) easily

OpenStudy (anonymous):

have no idea how the other answer came about, the answer is \(\{-2,1\}\)

OpenStudy (anonymous):

Yes your right and you set them equal to zero. But wouldnt that just be the x's? the -10 and the 11 would be the y's. SO your correct for the x's part.

OpenStudy (anonymous):

So**

OpenStudy (anonymous):

Yep -10 for 7*-2+4

OpenStudy (anonymous):

yeah you can find the second coordinate by replacing \(x\) by \(-2\) or by replacing \(x\) by \(1\) you get the same y because you are solving for where they are equal

OpenStudy (anonymous):

Thats all they are wanting. Thank you so much though for the expanding part I was so very confused. I will fan and medal you.